OpenStudy (hitaro9):

Finding equation of an Ellipse from 2 points?

OpenStudy (hitaro9):

Hey, so I'm given 2 points on an ellipse: (-1,1) and (.5/-2)

OpenStudy (hitaro9):

No idea what to do here.

OpenStudy (kc_kennylau):

Those points are the foci?

OpenStudy (hitaro9):

No, they're just points on the ellipse

OpenStudy (hitaro9):

Oh, and the ellipse is in standard position, so I'm assuming center is 0,0

OpenStudy (kc_kennylau):

But the two foci must be different, lest it be a circle...

OpenStudy (hitaro9):

Yeah.

OpenStudy (hitaro9):

I need to find the equation of the ellipse, not the foci

OpenStudy (hitaro9):

So I imagine finding the vertices would be easier and using x^2/a^2 + y^2/b^2 = 1 would be easier?

OpenStudy (kc_kennylau):

oh yeah i would use this formula

OpenStudy (hitaro9):

So how do I find a and b?

OpenStudy (kc_kennylau):

by plugging the two points in, i think?

OpenStudy (hitaro9):

Well that doesn't do me much good cause then I'm left with 2 variables?

OpenStudy (kc_kennylau):

but you have two equations?

OpenStudy (hitaro9):

Yeah, and if I set them equal I still have two variables, right?

OpenStudy (kc_kennylau):

Well you have two equations for two unknowns, so they're sufficient? :)

OpenStudy (hitaro9):

I dunno. I have: \[1/b^2 + 1/ c^2 =1 \] and \[.25/b^2 +4/c^2 =1\]

OpenStudy (hitaro9):

but I dunno what to do with them. If I set them equal I get \[.75/b^2 =3/c^2\] and that doesn't help me?

OpenStudy (kc_kennylau):

Don't set them equal :)

OpenStudy (hitaro9):

What do I do then?

OpenStudy (kc_kennylau):

Two ways to go: Substitution or Elimination :)

OpenStudy (hitaro9):

Ohhh

OpenStudy (hitaro9):

God I'm dumb, yeah, that would make sense

OpenStudy (kc_kennylau):

no you're not dumb :) your brain is just temporarily stuck :)

OpenStudy (hitaro9):

Appreciate the help man

OpenStudy (kc_kennylau):

no problem :)